The orbital speed of a planet about its star is 3.1
Answer #1

a) What is the magnitude of the force exerted by the star on the planet will be equal to centripetal force or m*v^2/r (m is mass of planet )

or (6*1024*9.61*108)/(1.6*1011) N or 36.0375 * 1021 N

B) What is the mass of the star?

now centripetal force must be equal to gravitational pull of the sun

or m1v^2/r = G*m1*m2/r^2 (m1 is mass of planet , m2 is mass of sun)

or m2 = v^2 *r /G

or m2 = (3.1*104)2*(1.6*1011)/6.67384